OpenStudy (anonymous):

Solving Rational Equations : 9/x+2=3/x 2+2x

OpenStudy (anonymous):

can you write that properly please using brackets. its really confusing. sorry.

OpenStudy (anonymous):

(9/x+2) = (3/x+2) +2x

OpenStudy (anonymous):

excellnt. so lets start by taking over the 3/(x+2) because we note that it has the same denom as the other side. so we get ... (9/x+2) - (3/x+2) = 2x so... from here we subtract the two ... 6/(x+2) = 2x ... now multiply the (x+2) to the other side which will result in ... 6 = 2x(x+2) ... expanding this we will get ... 6 = 2x^2 + 4x ... now bring the 6 over so we will get ... 0 = 2x^2 + 4x - 6 and now factoring out the two for easier solution we get ... 0 = 2 ( x^2 + 2x - 3) and this will factor out to.. 0 = 2(x+3)(x-1) soo ... from here your answer is x+3 = 0 ; hence x = -3 and x - 1 = 0 hence x = 1 So your final answers are x = -3 and 1.

OpenStudy (anonymous):

okay ! I think im getting the hang of it heh thanks. you are a greattt help :DD!

OpenStudy (anonymous):

that sounds good. haha npp :) check our khanacademy.com its a really super cool awesome resource for basic to super advnaced math and there are a lot of videos on it that teach about rational fractions. the videos are realyl quick and will help you out in terms of concepts

OpenStudy (anonymous):

thanks I'll be sure to check it out!